OpenStudy (anonymous):

Any Help please #CALCULUS #LIMITS Find lim(x->3)(√(x+1)-2)/(|-x^2-x+12|), can't solve it by myself

OpenStudy (luigi0210):

\[\LARGE \lim_{x \rightarrow 3} \frac{\sqrt{x+1}-2}{|-x^2-x+12|}\] Better?

OpenStudy (ranga):

I am getting different left and right limits. So the limit x->3 does not exist.

OpenStudy (ranga):

Find the roots of | -x^2 - x + 12 |. They are -4 and 3. -x^2 - x + 12 is positive when x = 3- and is negative when x = 3+. when x > 3+, | -x^2 - x + 12 | = x^2 + x - 12. limx->3+ { sqrt(x+1) - 2 } / { x^2 + x - 12 } = ? (apply L'H). Similar procedure for lim x->3-

OpenStudy (anonymous):

Thanks everyone for your fast responds and sorry for my slow reply(the school internet broke down again) The question is given by my lecturer, if it's non exist limit, may I know what should I give as an answer?

OpenStudy (ranga):

The limit does not exist because the left hand side limit is -1/28 and the right hand side limit is +1/28 and the two limits are not the same.

OpenStudy (anonymous):

Thank you, but when I multiply by conjugate, I still hey 0 over 0 when solving the limit, am I wrong at any point? Or if you don't mind, can you shows me some steps?

OpenStudy (ranga):

I wanted to get rid of the absolute bars in the denominator. To do that I found the roots so I can determine in what intervals the function is positive/negative. -x^2 - x + 12 = -(x^2 + x - 12) = -(x+4)(x-3) The roots are -4 and +3. -x^2 - x + 12 is negative in the intervals: (-infinity, -4) and (3, infinity) -x^2 - x + 12 is positive in the interval: (-4, 3). The point we are interested in is x = 3. Slightly to the left of 3, -x^2 - x + 12 is poistive. Slightly to the right of 3, -x^2 - x + 12 is negative. We can use this information to get rid of the absolute bars in the proper intervals.

OpenStudy (ranga):

x = 3-, -x^2 - x + 12 is positive and therefore, |-x^2 - x + 12| = -x^2 - x + 12 x = 3+, -x^2 - x + 12 is negative and therefore, |-x^2 - x + 12| = x^2 + x - 12

OpenStudy (ranga):

So the original limit of x->3 is split into two parts: lim x->3- and lim x->3+. lim x->3- { sqrt(x+1) - 2 } / { |-x^2 - x + 12| } = lim x->3- { sqrt(x+1) - 2 } / { -x^2 - x + 12 }. This is an indeterminate form 0/0. Therefore, we can apply L'H. Differentiate top and bottom separately: lim x->3- { 1/2 * (x+1)^(-1/2) } / { -2x - 1 } = { 1/2 * (4)^(-1/2) } / { -6 - 1 } = 1/2 * 1/2 * (-1/7) = -1/28. Can you do the right side limit?

OpenStudy (anonymous):

Thanks for being so kind to help, now I figure it out much better, I had spend few hours with it, thanks again!

OpenStudy (ranga):

You are welcome.